# Precompactness in matrix weighted Bourgain-Morrey spaces

### v1

@@ 2.2 Matrix weighted Bourgain-Morrey spaces
Lemma 2.14 (Proposition 3.2,[8])
For a matrix weight $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$, we have $0<w(x)\leq\|W(x)\|<\infty$ for a.e. $x\in{\mathbb{R}^{n}}$. Furthermore, $W$ satisfies a two weight, degenerate ellipticity condition: for $\xi\in\mathbb{R}^{d}$,
$w(x)|\xi|^{p}\leq|W^{1/p}(x)\xi|^{p}\leq\|W(x)\||\xi|^{p}.$ (2.7)
Remark 2.15
In [8, Proposition 3.2], it is assumed that $\xi\in\mathbb{R}^{d}$ but it also works for $\xi\in\mathbb{C}^{d}$.
Proposition 3
@@
  • Proof
    Clearly, if $\vec{f}(x)=\vec{g}(x)$ a.e., then $\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}=0$. Then we apply Lemma 2.14 to prove the converse. By the degenerate ellipticity condition (2.7), we have
    00 0 $=\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}$
    $\geq\left\|\left\{W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|\vec{f}(y)-\vec{g}(y)|^{p}w(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}.$
    Since $w(y)>0$ a.e., it follows that $\vec{f}(y)-\vec{g}(y)=0$ a.e.. $\Box$
@@ 3 Sufficient conditions for precompact sets
Definition 3.1
(i) Suppose that $(X,\rho)$ is a metric space. Let $A$ be a set of $X$ and let $\varepsilon>0$. A set $E\subset X$ is called an $\varepsilon$-net for $A$ if every point $a\in A$, there exists a point $e\in E$ such that $\rho(a,e)<\varepsilon$.
(ii) A set $A$ is called totally bounded if, for each $\varepsilon>0$, it possesses a finite $\varepsilon$-net.
@@
  • Proof of Theorem 3.2.
    Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i) and (ii). Given $\varepsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\varepsilon$-net of $\mathcal{F}$. Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$. Then from condition (i), there exist a positive integer $m$ large enough such that
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (3.2)
@@
    $\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R_{a}}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (3.3)
    There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\cup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
    For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
    $\varPhi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
@@
    $\left|W(x)\left(\vec{f}(x)-\vec{f}_{Q_{j}}\right)\right|\chi_{Q_{j}}(x)$ $=\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}W(x)\left(\vec{f}(x)-\vec{f}(y)\right){\mathrm{d}}y\right|\chi_{Q_{j}}(x)$ (3.4)
    $\leq\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x).$ (3.5)
    We split $\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}$ into three parts:
    $\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    +$\quad+\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    +$\quad+\left\{\sum_{k<-m}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
@@
    We first estimate $S_{1}$. By (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we have
    $S_{1}$$\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.6)
    ×$\quad\times\left.\left(\int_{Q}\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.7)
    $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.8)
    ×$\quad\times\left.\left(\frac{1}{|Q_{j}|}\int_{Q_{j}}\int_{Q}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$ (3.9)
    =$=\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.10)
    ×$\quad\times\left.\left(\frac{1}{|Q_{j}|}\int_{Q}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.11)
    $=2^{-an/p}\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.12)
    ×$\quad\times\left.\left(\int_{Q}\int_{x-Q_{j}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.13)
    where $x-Q_{j}:=\{x-y:y\in Q_{j}\}$. Since $x\in Q\subset Q_{j}$ for some $j\in\{1,2,\ldots,N\}$, we have $x-Q_{j}\subset R_{a}$. Hence by (3.3), we obtain
@@
    ×$\quad\left.\times\left(\int_{Q}\left|W(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    $\leq\varepsilon.$
    As for $S_{2}$, for each $Q\in{\mathcal{D}}_{k}$ where $k=-m,-m+1,\ldots,-a$, there are $2^{(-k-a)n}$ cubes $Q_{j}\in{\mathcal{D}}_{-a}$ such that $\cup Q_{j}=Q$. Denote by $Q_{j_{\ell}}$, $\ell=1,2,\ldots,2^{(-k-a)n}$ these cubes. Then by (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we obtain
    $S_{2}$ =$=\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.14)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j_{\ell}}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.15)
    $\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.16)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.17)
    $\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.18)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$ (3.19)
    $=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.20)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.21)
    $=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\right.$ (3.22)
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{x-Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (3.23)
    where $x-Q_{j_{\ell}}:=\{x-y:y\in Q_{j_{\ell}}\}$. Note that $x-Q_{j_{\ell}}\subset R_{a}$ when $x\in Q_{j_{\ell}}$. By (3.3), we have
@@
    ×$\quad\times\left.\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    $\leq\varepsilon.$
@@
    $S_{3}$ =$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\int_{Q_{\ell}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
    =$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\sum_{j=1}^{2^{(m-a)n}}\int_{Q_{j}}\left|W(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}.$
    Similarly as $S_{2}$, from (3.4), Jensen’s inequality ($p\geq 1$), the Fubini theorem, and (3.3), we get that $S_{3}\leq\varepsilon$. Together with the estimates of $S_{1}$, $S_{2}$, $S_{3}$, we obtain
    $\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W(x)(\vec{f}(x)\chi_{R_{m}}-\varPhi(\vec{f})(x))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}\leq 3\varepsilon.$ (3.24)
@@
    $\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$
    $=\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}.$
    Thus via (3.2) and (3.24), we have
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq 4\varepsilon.$ (3.25)
    From (3.25), it suffices to show that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$.
@@
    $\sup_{\vec{f}\in\mathcal{F}}\|\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq\sup_{\vec{f}\in\mathcal{F}}\|\varPhi(\vec{f})-\vec{f}\|_{M_{p}^{t,r}(W)}+\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
@@
    $|W^{1/p}(x)\varPhi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (3.26)
@@
    $|\varPhi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (3.27)
    which implies $|\vec{f}_{Q_{j}}|<\infty$, $j=1,2,\ldots,N$. From this and the entries of $W$ is locally integrable, we see that $\varPhi$ is a map from $\mathcal{F}$ to $\mathcal{B}$, a finite dimensional Banach subspace of $M_{p}^{t,r}(W)$. Note that $\varPhi(\mathcal{F})\subset\mathcal{B}$ is bounded, and hence is totally bounded. The proof of Theorem 3.2 is complete. $\Box$
@@
is the norm of the scalar weighted Bourgain-Morrey space in (2.6).
Lemma 3.5 (p. 262, [33])
@@
  • Proof
    Note that $\mathcal{F}\subset\mathcal{W}^{1,p,t,r}(W)$ satisfies conditions (i)-(ii) if and only if $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ satisfies conditions (i)-(ii) of Theorem 3.2 ($d=1$) and :=$\nabla\mathcal{F}:=\{\nabla f:f\in\mathcal{F}\}\subset M^{t,r}_{p}(W)$ satisfies (i)-(ii) of Theorem 3.2 ($d=n$). Hence by Theorem 3.2, we obtain that both $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ and $\nabla\mathcal{F}\subset M^{t,r}_{p}(W)$ are totally bounded. Then Corollary 3.6 follows from Lemma 3.5. $\Box$

### v2

@@ 2.2 Matrix weighted Bourgain-Morrey spaces
Lemma 2.17 (Proposition 3.2,[10]).
Let $1\leq p<\infty$. For a matrix weight $W:{\mathbb{R}^{n}}\to M_{d}(\mathbb{C})$, we have $0<w(x)\leq\|W(x)\|<\infty$ for a.e. $x\in{\mathbb{R}^{n}}$. Furthermore, $W$ satisfies a two weight, degenerate ellipticity condition: for $\xi\in\mathbb{R}^{d}$,
$w(x)|\xi|^{p}\leq|W^{1/p}(x)\xi|^{p}\leq\|W(x)\||\xi|^{p}.$ (7)
Remark 2.18.
In [10, Proposition 3.2], it is assumed that $\xi\in\mathbb{R}^{d}$ but it also works for $\xi\in\mathbb{C}^{d}$.
Proposition 2.19.
@@
Proof.
Clearly, if $\vec{f}(x)=\vec{g}(x)$ a.e., then $\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}=0$. Then we apply Lemma 2.17 to prove the converse. By the degenerate ellipticity condition (7), we have
$0=\|\vec{f}-\vec{g}\|_{M_{p}^{t,r}(W)}\geq\left\|\left\{W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|\vec{f}(y)-\vec{g}(y)|^{p}w(y){\mathrm{d}}y\right)^{1/p}\right\}_{j\in\mathbb{Z},k\in\mathbb{Z}^{n}}\right\|_{\ell^{r}}.$
Since $w(y)>0$ a.e., it follows that $\vec{f}(y)-\vec{g}(y)=0$ a.e.. ∎
@@ 3 Sufficient conditions for precompact sets
Definition 3.1.
(i) Suppose that $(X,\rho)$ is a metric space. Let $A$ be a set of $X$ and let $\epsilon>0$. A set $E\subset X$ is called an $\epsilon$-net for $A$ if every point $a\in A$, there exists a point $e\in E$ such that $\rho(a,e)<\epsilon$.
(ii) A set $A$ is called totally bounded if, for each $\epsilon>0$, it possesses a finite $\epsilon$-net.
@@
Now we begin to show Theorem 3.2.
Proof.
Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i) and (ii). Given $\epsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\epsilon$-net of $\mathcal{F}$. Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$. Then from condition (i), there exist a positive integer $m$ large enough such that
$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (9)
@@
$\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R_{a}}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (10)
There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\bigcup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$. For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
$\Phi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
@@
$\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}_{Q_{j}}\right)\right|\chi_{Q_{j}}(x)$ $=\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right){\mathrm{d}}y\right|\chi_{Q_{j}}(x)$
$\leq\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x).$ (11)
We split $\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}$ into three parts:
$\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$\quad+\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$\quad+\left\{\sum_{k<-m}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
@@
We first estimate $S_{1}$. By (3), Jensen’s inequality ($p\geq 1$), the Fubini theorem, we have
$S_{1}$ $\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|\frac{1}{|Q_{j}|}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\chi_{Q_{j}}(x)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$\leq\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{j}|}\int_{Q_{j}}\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$
$=\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{j}|}\int_{Q}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$=2^{-an/p}\left\{\sum_{k>-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\int_{Q}\int_{x-Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (12)
where $x-Q_{j}:=\{x-y:y\in Q_{j}\}$. Since $x\in Q\subset Q_{j}$ for some $j\in\{1,2,\ldots,N\}$, we have $x-Q_{j}\subset R_{a}$. Hence by (10), we obtain
@@
As for $S_{2}$, for each $Q\in{\mathcal{D}}_{k}$ where $k=-m,-m+1,\ldots,-a$, there are $2^{(-k-a)n}$ cubes $Q_{j}\in{\mathcal{D}}_{-a}$ such that $\cup Q_{j}=Q$. Denote by $Q_{j_{\ell}}$, $\ell=1,2,\ldots,2^{(-k-a)n}$ these cubes. Then by (3), Jensen’s inequality ($p\geq 1$), and the Fubini theorem, we obtain
$S_{2}$ $=\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j_{\ell}}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\left|\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|{\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$\leq\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\frac{1}{|Q_{j_{\ell}}|}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right\}^{1/r}$
$=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$=2^{-an/p}\left\{\sum_{k=-m}^{-a}\sum_{Q\in\mathcal{D}_{k},Q\subset R_{m}}W(Q)^{r/t-r/p}\left(\sum_{\ell=1}^{2^{(-k-a)n}}\int_{Q_{j_{\ell}}}\int_{x-Q_{j_{\ell}}}\left|W^{1/p}(x)\left(\vec{f}(x)-\vec{f}(x-y)\right)\right|^{p}{\mathrm{d}}y{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$ (13)
where $x-Q_{j_{\ell}}:=\{x-y:y\in Q_{j_{\ell}}\}$. Note that $x-Q_{j_{\ell}}\subset R_{a}$ when $x\in Q_{j_{\ell}}$. By (10), we have
@@
$S_{3}$ $=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\int_{Q_{\ell}}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x)\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}$
$=\left\{\sum_{k<-m}\sum_{\ell=1}^{2^{n}}W(Q_{\ell})^{r/t-r/p}\left(\sum_{j=1}^{2^{(m-a)n}}\int_{Q_{j}}\left|W^{1/p}(x)\left(\vec{f}(x)\chi_{R_{m}}-\vec{f}_{Q_{j}}\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}.$
Similarly as $S_{2}$, from (3), Jensen’s inequality ($p\geq 1$), the Fubini theorem, and (10), we get that $S_{3}\leq\epsilon$. Together with the estimates of $S_{1}$, $S_{2}$, $S_{3}$, we obtain
$\left\{\sum_{k\in\mathbb{Z}}\sum_{Q\in\mathcal{D}_{k}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)(\vec{f}(x)\chi_{R_{m}}-\Phi(\vec{f})(x))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right\}^{1/r}\leq 3\epsilon.$ (14)
@@
$\|\vec{f}-\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}=\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}.$
Thus via (9) and (14), we have
$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq 4\epsilon.$ (15)
From (15), it suffices to show that $\Phi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$.
@@
$\sup_{\vec{f}\in\mathcal{F}}\|\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}\leq\sup_{\vec{f}\in\mathcal{F}}\|\Phi(\vec{f})-\vec{f}\|_{M_{p}^{t,r}(W)}+\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
@@
$|W^{1/p}(x)\Phi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (16)
@@
$|\Phi(\vec{f})(x)|<\infty\quad{\rm a.e.}\quad x\in{\mathbb{R}^{n}}.$ (17)
which implies $|\vec{f}_{Q_{j}}|<\infty$, $j=1,2,\ldots,N$. From this and the entries of $W$ is locally integrable, we see that $\Phi$ is a map from $\mathcal{F}$ to $\mathcal{B}$, a finite dimensional Banach subspace of $M_{p}^{t,r}(W)$. Note that $\Phi(\mathcal{F})\subset\mathcal{B}$ is bounded, and hence is totally bounded. The proof of Theorem 3.2 is complete. ∎
@@
is the norm of the scalar weighted Bourgain-Morrey space in (6).
Lemma 3.5 (p. 262, [36]).
@@
Proof.
Note that $\mathcal{F}\subset\mathcal{W}^{1,p,t,r}(W)$ satisfies conditions (i)-(ii) if and only if $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ satisfies conditions (i)-(ii) of Theorem 3.2 ($d=1$) and $\nabla\mathcal{F}:=\{\nabla f:f\in\mathcal{F}\}\subset M^{t,r}_{p}(W)$ satisfies (i)-(ii) of Theorem 3.2 ($d=n$). Hence by Theorem 3.2, we obtain that both $\mathcal{F}\subset M^{t,r}_{p}(\omega)$ and $\nabla\mathcal{F}\subset M^{t,r}_{p}(W)$ are totally bounded. Then Corollary 3.6 follows from Lemma 3.5. ∎